Paragraph number one is simply plain text. It has several sentences. Three, in fact, although the final one keeps going and going and going for a long time because I wanted to make sure that the third paragraph would span several lines in the final document.

The second paragraph is separated from the first paragraph by a single blank line. That’s all it takes to create paragraphs: a single blank line between the input text.